Here is the problem I'm stuck on...
A commercial jet travels 1,620 miles in the same amount of time that a corporate jet travels 1,260 miles. The rate of the commercial jet is 120 miles per hour faster than the rate of the corporate jet. Find the rate of each jet.

Answer by mananth(16949)

Corporate jet B x mph
Commercial jet A 120 faster

Corporate jet 1260 miles
Commercial jet 1620 miles
Corporate jet time = Commercial jet time

1260 / x = 1620 /( x + 120
cross multiply
1260 ( x + 120 ) = 1620 x
1260 x + 151200 = 1620 x
1620 x - 1260 x = 151200
360 x = 151200
/ 360
x = 420 mph
420 mph Corporate jet
540 mph Commercial jet